# Classify all representation which contains a Steinberg in its hyperspecial subgroup

### 24 Mar 2026 v1

@@ Abstract
This paper addresses Question 1 posed by Dipendra Prasad in his recent problem list: classify all irreducible smooth representations $\pi$ of an unramified reductive $p$-adic group $G$ (over a characteristic zero local field) such that the space of $K_{+}$-fixed vectors $V^{K_{+}}$ (viewed as a representation of the finite reductive group $K/K_{+}$) contains the Steinberg representation. We prove that any such representation $\pi$ must be Iwahori-spherical, hence a subquotient of some unramified principal series $I(\chi)=\operatorname{Ind}_{B}^{G}\chi$. By a detailed analysis of the action of the Iwahori–Hecke algebra on the Iwahori-fixed space $V^{I}$, we show that in each principal series $I(\chi)$ there exists exactly one irreducible subquotient containing Steinberg, and we give an explicit construction of this subquotient. Hence this can give us a bijection between W-orbits of unramified characters and irreducible representations containing Steinberg in its hyperspecial subgroup. This classification is richer than Prasad’s expectation: besides twists of the Steinberg representation and generic unramified representations, every principal series contributes a unique subquotient with the required property. The proof relies on establishing an isomorphism between the finite Iwahori–Hecke algebra of the finite Lie group $K/K_{+}$ and the finite Hecke algebra $H(K,I)$ of the $p$-adic group (compactly supported $I$-bi-invariant functions on $K$), and comparing their actions on $V^{I}$.
Keywords: $p$-adic groups, unramified reductive groups, Steinberg representation, Iwahori subgroup, Hecke algebras, unramified principal series, Moy–Prasad filtration.
@@ 1 Introduction
Prasad suggested that such representations should fall into two families: (1) twists of the Steinberg representation of $G(F)$; (2) generic, unramified representations (i.e., irreducible spherical principal series). However, our investigation reveals a more nuanced picture: every unramified principal series $I(\chi)=\operatorname{Ind}_{B}^{G}\chi$ (where $B$ is a Borel subgroup and $\chi$ an unramified character of a maximal torus) contains exactly one irreducible subquotient that contains the Steinberg representation. In particular, the class of representations satisfying Prasad’s condition is strictly larger than his expectation; each principal series contributes a unique such subquotient, hence this can give us a bijection between W-orbits of unramified characters and irreducible representations containing Steinberg in its hyperspecial subgroup.
The main steps of our argument are as follows. Let $(\pi,V)$ be an irreducible smooth representation of $G(F)$ with $V^{K_{+}}$ containing the Steinberg representation. By the depth-zero theory of Moy–Prasad [8, 9], $\pi$ must be of depth zero. Using Lemma 3.2 and the Borel–Casselman theorem [1], we deduce that $\pi$ is Iwahori-spherical, i.e., $V^{I}\neq 0$ for an Iwahori subgroup $I\subset K$. Hence $\pi$ is a subquotient of some unramified principal series $I(\chi)$. The space $V^{K_{+}}$ is a representation of the finite reductive group $K/K_{+}$; by Lemma 3.1 all its irreducible constituents lie in the principal Harish-Chandra series, i.e., are subrepresentations of $\operatorname{Ind}_{I/K_{+}}^{K/K_{+}}\mathbf{1}$. Taking $I/K_{+}$-fixed points yields $V^{I}$, which carries two natural actions: one from the Hecke algebra $H_{\mathrm{fin}}=\operatorname{End}_{K/K_{+}}(\operatorname{Ind}_{I/K_{+}}^{K/K_{+}}\mathbf{1})$ (coming from finite group theory), and one from the finite Hecke algebra $H(K,I)$ of the $p$-adic group (compactly supported $I$-bi-invariant functions on $K$). In Section 6 we prove that these two actions coincide on $V^{I}$ (Lemma 6.2), allowing us to transfer known results about the Steinberg representation from finite groups to the $p$-adic setting.
The multiplicity of the Steinberg representation in $V^{K_{+}}$ equals the dimension of the space
$W=\{v\in V^{I}\mid T_{\alpha}v=-v\text{ for\ every \ simple \ root \ }\alpha\},$
where $T_{\alpha}$ are the standard generators of the Hecke algebra. Using Casselman’s basis $\{\phi_{w,\chi}\}$ of $I(\chi)^{I}$ and Lansky’s formulas for the action of affine reflections [7] (adapted to unnormalized induction), we solve this eigenvalue problem in $I(\chi)^{I}$. Lemma 7.2 shows that the solution space is one‑dimensional; a non‑zero vector is explicitly given by
@@
Consequently $\dim W=1$, so the Steinberg representation occurs with multiplicity one in $V^{K_{+}}$. Moreover, letting $V$ be the subrepresentation of $I(\chi)$ generated by $v_{0}$ and $V_{0}$ the maximal subrepresentation of $V$ not containing $v_{0}$, the quotient $V/V_{0}$ is the unique irreducible subquotient of $I(\chi)$ that contains Steinberg. This completes the classification.
The paper is organized as follows. Section 2 recalls the Moy–Prasad filtration and unrefined minimal $K$-types. Section 3 analyses the cuspidal support of the irreducible components of $V^{K_{+}}$ and proves that a representation containing Steinberg must be Iwahori-spherical. Section 4 collects facts about unramified principal series: Casselman’s basis and Lansky’s action of affine reflections. Section 5 reviews the representation theory of finite Lie groups, focusing on the principal series and the associated Iwahori–Hecke algebra. Section 6 establishes the isomorphism between the finite Hecke algebra and the $p$-adic finite Hecke algebra $H(K,I)$, and proves that their actions on $V^{I}$ coincide. Section 7 contains the core lemma and the proof of the main theorem.
@@ 2.1 Moy–Prasad filtrations
For a point $x$ in the Bruhat–Tits building of $G/k$, let $\mathcal{P}_{x}=\mathcal{P}_{x}\cap G(k)$ be the corresponding parahoric subgroup of $G(k)$. Moy and Prasad associate to $x$ a decreasing filtration $\{\mathcal{P}_{x,r}\}_{r\geq 0}$ of $\mathcal{P}_{x}$ by open compact subgroups, defined as follows: for a non-negative real number $r$, let $\mathcal{P}_{x,r}$ be the subgroup generated by certain congruence subgroups of the maximal torus and by the root subgroups $U_{\psi}$ for all affine roots $\psi$ satisfying $\psi(x)\geq r$; see [8, 2.6] for the precise definition. One sets $\mathcal{P}_{x,0}=\mathcal{P}_{x}$ and $\mathcal{P}_{x,r+}=\bigcup_{s>r}\mathcal{P}_{x,s}$. For $r>0$ the quotients $\mathcal{P}_{x,r}/\mathcal{P}_{x,r+}$ are abelian.
There is an associated filtration of the Lie algebra $\mathfrak{g}$. Let $\mathfrak{g}_{x,r}$ be the $\mathfrak{o}$-lattice in $\mathfrak{g}$ obtained from the smooth $\mathfrak{o}$-group scheme corresponding to $\mathcal{P}_{x}$; its definition is given in [9, 3.2]. Similarly, one defines a filtration $\{\mathfrak{g}_{x,r}^{*}\}_{r\in\mathbb{R}}$ of the dual $\mathfrak{g}^{*}$ by
@@
is non-degenerate and $\mathbf{M}_{x}(\mathfrak{f})$-equivariant, where $\mathbf{M}_{x}$ is the reductive quotient of the special fibre of the integral model attached to $x$; cf. [9, 3.7].
For a character $\chi$ of $\mathcal{P}_{x,r}$ trivial on $\mathcal{P}_{x,r+}$ (with $r>0$), one can realize $\chi$ as a character of the abelian group $\mathcal{P}_{x,r}/\mathcal{P}_{x,r+}$. Via the isomorphism $\mathcal{P}_{x,r}/\mathcal{P}_{x,r+}\cong\mathfrak{g}_{x,r}/\mathfrak{g}_{x,r+}$ of [9, 3.8] and the above duality, such a character corresponds to a coset
@@ 3 The cuspidal support of the irreducible components of $K/K_{+}$
For a fixed Iwahori subgroup $I$ of $G$ , By [6, Theorem 8.4.10(2)], we know that $I/I_{+}$ is the minimal Levi of $K/K_{+}$. Since every finite Lie group is quasi-split, $I/I_{+}$ is a torus, and $I/K_{+}$ is the Borel containing this torus.
@@ 4.2 Effect of affine reflections on the Iwahori-fixed vectors
From the results of [7, Theorem 3.1] one deduces the action of the characteristic functions of the double cosets IsIIsI ($s\in S_{\mathrm{aff}}$) in the Iwahori–Hecke algebra $\mathcal{H}(G,I)$. The following corollary records the formulas we need, adapted to unnormalized induction.

### 31 Mar 2026 v2

@@ Abstract
This paper addresses Question 1 posed by Dipendra Prasad in his recent problem list [15]: classify all irreducible smooth representations $\pi$ of an unramified reductive $p$-adic group $G$ (over a characteristic zero local field) such that the space of $K_{+}$-fixed vectors $V^{K_{+}}$ (viewed as a representation of the finite reductive group $K/K_{+}$) contains the Steinberg representation. Prasad initially expected that such representations should be twists of the Steinberg representation or generic unramified representations. In a subsequent private communication, he refined the expectation: they should be precisely the unique generic irreducible subquotients of unramified principal series. We prove that any such $\pi$ must be Iwahori-spherical, hence a subquotient of some unramified principal series $I(\chi)=\operatorname{Ind}_{B}^{G}\chi$. By a detailed analysis of the Iwahori-Hecke algebra action on the Iwahori-fixed space $V^{I}$, we show that each principal series $I(\chi)$ contains exactly one irreducible subquotient whose $K_{+}$-fixed vectors contain the Steinberg representation, and we give an explicit construction of this subquotient. Consequently, we obtain a bijection between $W$-orbits of unramified characters and irreducible representations containing the Steinberg representation in their hyperspecial subgroup. The proof is self-contained and uses only elementary Hecke algebra methods. In particular, the reduction from the $K$-action on $K_{+}$-fixed vectors to modules over the finite Hecke algebra $H(K,I)$ is spelled out in detail, a step that has not been made explicit in the literature.
@@ 1 Introduction
Prasad initially suggested that such representations should fall into two families: (1) twists of the Steinberg representation of $G(F)$; (2) generic unramified representations (i.e., irreducible spherical principal series). In a subsequent private communication, he refined this expectation: the representations in question should be precisely the unique generic irreducible subquotients of unramified principal series.
The present paper confirms this refined expectation and gives a complete classification. Our main results are as follows.
  • Existence and uniqueness. For every unramified principal series $I(\chi)=\operatorname{Ind}_{B}^{G}\chi$ (where $B$ is a Borel subgroup and $\chi$ an unramified character of a maximal torus), there exists exactly one irreducible subquotient whose $K_{+}$-fixed vectors contain the Steinberg representation. We give an explicit construction of this subquotient via the Iwahori-Hecke algebra.
  • Parametrization. The construction yields a bijection between $W$-orbits of unramified characters (where $W$ is the relative Weyl group) and irreducible representations of $G(F)$ whose $K_{+}$-fixed vectors contain the Steinberg representation.
Our proof is self-contained and uses only elementary methods from the theory of Iwahori-Hecke algebras. The main steps are:
  1. 1.
    Show that any representation satisfying Prasad’s condition must be Iwahori-spherical (hence a subquotient of some $I(\chi)$) using the Moy–Prasad theory of depth-zero representations and the Borel–Casselman theorem.
  2. 2.
    Reduce the problem to an eigenvalue problem on the Iwahori-fixed space $V^{I}$ by comparing two actions on $V^{I}$: one coming from the finite Hecke algebra $H(K,I)$ of the $p$-adic group, and the other from the Iwahori–Hecke algebra of the finite Lie group $K/K_{+}$. We prove that these two actions coincide (Lemma 6.2), a step that has not been spelled out clearly in the literature.
  3. 3.
    Solve the eigenvalue problem using Casselman’s basis $\{\phi_{w,\chi}\}$ of $I(\chi)^{I}$ and Lansky’s formulas for the action of affine reflections. The unique (up to scalar) solution is given explicitly by
@@
  4. 4.
    Show that the subrepresentation generated by $v_{0}$ contains a unique irreducible subquotient with the desired property. And we also show that this subquotient is the unique generic subquotient in a unramified principal series.
Relation to previous work
After the first version of this paper appeared on arXiv, we were informed by Manish Mishra and Maarten Solleveld that related results exist in the literature. We are very grateful to them for their careful reading and for pointing out the connections.
  • Manish Mishra [9] has studied generic representations in $L$-packets. In Theorem 12 of [9], he proved the existence and uniqueness theorem under the framework of unramified L-packets. In the split case, Theorem 3 of his joint work with Basudev Pattanayak [10] gives an explicit Hecke-algebra model for the unique generic constituent of an unramified principal series. These results, which are part of a more general framework for $L$ packages, imply our classification in a broader context.
  • Maarten Solleveld [16] developed a comprehensive theory of principal series representations for quasi-split reductive $p$-adic groups. In Theorem 3.4 of [16], he established that the condition used by Prasad is equivalent to the representation being generic.(To see that one has to know as well the reduction steps from the K-action on K+-invariant vector to modules over H(K,I), which is contained in this article. ) His work, which is much more general, provides a conceptual explanation for our results.
Professor Solleveld also kindly pointed out that the reduction steps from the $K$-action on $K_{+}$-fixed vectors to modules over H(K,I)H(K,I)—which we work out in detail—are not explicitly written elsewhere, and that this part of our work can be regarded as an original contribution. We are deeply indebted to them for their generosity and for helping us locate our work within the larger framework.
Original contributions of this paper
While the existence and uniqueness of generic subquotients of unramified principal series are known from general theory, our work provides the following new elements:
  1. 1.
    Explicit construction. We give an explicit formula for a basis vector $v_{0}$ in the Iwahori-fixed space that generates the desired subquotient. This explicit description is new and may be useful for further computations .
  2. 2.
    Self-contained elementary proof. Our proof uses only Hecke algebra calculations and does not rely on the more advanced machinery of $L$-packets or the local Langlands correspondence.
  3. 3.
    Detailed reduction steps. The comparison between the finite Hecke algebra $H(K,I)$ and the Iwahori–Hecke algebra of the finite Lie group $K/K_{+}$ (Lemma 6.2) is made completely explicit. As noted by Solleveld, this step, has not been written out clearly in the literature.
Organization of the paper
Section 2 recalls the Moy–Prasad filtration and the theory of unrefined minimal $K$-types. Section 3 analyzes the cuspidal support of the irreducible components of $V^{K_{+}}$ and proves that any representation containing the Steinberg representation must be Iwahori-spherical. Section 4 collects necessary facts about unramified principal series, including Casselman’s basis and Lansky’s formulas for the action of affine reflections. Section 5 reviews the representation theory of finite Lie groups, focusing on the principal series and the associated Iwahori–Hecke algebra. Section 6 establishes an isomorphism between the finite Hecke algebra of the finite group and the finite Hecke algebra $H(K,I)$ of the $p$-adic group, and proves that their actions on $V^{I}$ coincide. Section 7 contains the core lemma and the proof of the main theorem, including the explicit construction of the subquotient. Finally, we include a corollary giving a bijection between $W$-orbits of unramified characters and the classified representations. Section 8 provides readers with another view of this result.
@@ 2.1 Moy–Prasad filtrations
For a point $x$ in the Bruhat–Tits building of $G/k$, let $\mathcal{P}_{x}=\mathcal{P}_{x}\cap G(k)$ be the corresponding parahoric subgroup of $G(k)$. Moy and Prasad associate to $x$ a decreasing filtration $\{\mathcal{P}_{x,r}\}_{r\geq 0}$ of $\mathcal{P}_{x}$ by open compact subgroups, defined as follows: for a non-negative real number $r$, let $\mathcal{P}_{x,r}$ be the subgroup generated by certain congruence subgroups of the maximal torus and by the root subgroups $U_{\psi}$ for all affine roots $\psi$ satisfying $\psi(x)\geq r$; see [11, 2.6] for the precise definition. One sets $\mathcal{P}_{x,0}=\mathcal{P}_{x}$ and $\mathcal{P}_{x,r+}=\bigcup_{s>r}\mathcal{P}_{x,s}$. For $r>0$ the quotients $\mathcal{P}_{x,r}/\mathcal{P}_{x,r+}$ are abelian.
There is an associated filtration of the Lie algebra $\mathfrak{g}$. Let $\mathfrak{g}_{x,r}$ be the $\mathfrak{o}$-lattice in $\mathfrak{g}$ obtained from the smooth $\mathfrak{o}$-group scheme corresponding to $\mathcal{P}_{x}$; its definition is given in [12, 3.2]. Similarly, one defines a filtration $\{\mathfrak{g}_{x,r}^{*}\}_{r\in\mathbb{R}}$ of the dual $\mathfrak{g}^{*}$ by
@@
is non-degenerate and $\mathbf{M}_{x}(\mathfrak{f})$-equivariant, where $\mathbf{M}_{x}$ is the reductive quotient of the special fibre of the integral model attached to $x$; cf. [12, 3.7].
For a character $\chi$ of $\mathcal{P}_{x,r}$ trivial on $\mathcal{P}_{x,r+}$ (with $r>0$), one can realize $\chi$ as a character of the abelian group $\mathcal{P}_{x,r}/\mathcal{P}_{x,r+}$. Via the isomorphism $\mathcal{P}_{x,r}/\mathcal{P}_{x,r+}\cong\mathfrak{g}_{x,r}/\mathfrak{g}_{x,r+}$ of [12, 3.8] and the above duality, such a character corresponds to a coset
@@ 3 The cuspidal support of the irreducible components of $K/K_{+}$
For a fixed Iwahori subgroup $I$ of $G$ , By [7, Theorem 8.4.10(2)], we know that $I/I_{+}$ is the minimal Levi of $K/K_{+}$. Since every finite Lie group is quasi-split, $I/I_{+}$ is a torus, and $I/K_{+}$ is the Borel containing this torus.
@@ 4.2 Effect of affine reflections on the Iwahori-fixed vectors
From the results of [8, Theorem 3.1] one deduces the action of the characteristic functions of the double cosets IsIIsI ($s\in S_{\mathrm{aff}}$) in the Iwahori–Hecke algebra $\mathcal{H}(G,I)$. The following corollary records the formulas we need, adapted to unnormalized induction.
@@ 7.1 Main Theorem
Remark: In Theorem 3.4 of [16], Professor Solleveld established that the condition used by Prasad is equivalent to the representation being generic.(To see that one has to know as well the reduction steps from the K-action on K+-invariant vector to modules over H(K,I), which is contained in this article. )
8 Another Perspective: generic representation
Theorem 8.1.
The representation constructed above is the unique generic subquotient in the unramified principal series $I(\chi)$.
Proof.
It is well known that there is exactly one subquotient of $I(\chi)$ is generic. And the work of Moeglin and Waldspurger [13] told us an irreducible representation $(\pi,V)$ is generic if and only if the Wavefront set of $\pi$ is the regular orbit. From Example 2.3.8 of Okada’s paper [14] we can see that if $V^{K_{+}}$ contains Steinberg, then the geometry wavefront set of this representation must be the regular nilpotent orbit. Hence the representation that we construct above is the unique generic subquotient. ∎